# Computational Complexity of Sufficiency in Decision Problems

### v2

@@ 4 Encoding-Regime Separation
This encoding-regime separation explains why some domains admit tractable model selection under explicit-state assumptions, while other domains (or encodings) exhibit worst-case hardness that forces heuristic approaches.
@@ 5 Tractable Special Cases: When You Can Solve It
We distinguish the encodings of Section 2.4. The tractability results below state the model assumption explicitly.
@@ 5.4 Practical Implications
6 Implications for Practice: Why Over-Modeling Is Optimal
This section states corollaries for engineering practice. Within the formal model, the complexity results of Sections 3 and 4 shift parts of this workflow from informal judgment toward explicit, checkable criteria. The observed behaviors—configuration over-specification, absence of automated minimization tools, heuristic model selection—are not failures of discipline but provably optimal responses to computational constraints under the stated cost model.
Some common prescriptions implicitly require exact minimization of sufficient parameter sets. In the worst case, that task is coNP-complete in our model, so we should calibrate critiques and expectations accordingly. A rational response is to include everything and pay linear maintenance costs, rather than attempt exponential minimization costs.
@@ 6.1 Configuration Simplification is SUFFICIENCY-CHECK
Real engineering problems reduce directly to the decision problems studied in this paper.
Theorem 6.1 (Configuration Simplification Reduces to SUFFICIENCY-CHECK).
Given a software system with configuration parameters $P=\{p_{1},\ldots,p_{n}\}$ and observed behaviors $B=\{b_{1},\ldots,b_{m}\}$, the problem of determining whether parameter subset $I\subseteq P$ preserves all behaviors is equivalent to SUFFICIENCY-CHECK.
@@
  • Actions $A=B$ (each behavior is an action)
@@
  • Utility $U(b,s)=1$ if behavior $b$ occurs under configuration $s$, else $U(b,s)=0$
Then $\text{Opt}(s)=\{b\in B:b\text{ occurs under configuration }s\}$.
@@
Remark 6.2.
This reduction is parsimonious: configuration simplification can be cast as an instance of SUFFICIENCY-CHECK under a direct encoding.
6.2 Computational Rationality of Over-Modeling
We now prove that over-specification is an optimal engineering strategy given the stated cost model and complexity constraints.
Theorem 6.3 (Rational Over-Modeling).
@@ 6.2 Computational Rationality of Over-Modeling
When SUFFICIENCY-CHECK is coNP-complete (Theorem 3.6):
@@
    Naive exhaustive search inspects $2^{n}$ candidate subsets, so the brute-force finding cost grows exponentially in $n$; more formally, exhaustive search yields an upper bound $C_{\text{find}}(n)=O(2^{n})$. A matching *unconditional* lower bound of $2^{\Omega(n)}$ does not follow from coNP-completeness alone. Instead, coNP-completeness implies there is no polynomial‑time algorithm for arbitrary inputs unless $\textsf{P}=\textsf{coNP}$. Stronger assumptions (e.g., the Exponential Time Hypothesis) yield explicit worst-case lower bounds: under ETH, SUFFICIENCY-CHECK has a $2^{\Omega(n)}$ worst-case time lower bound in the succinct encoding (witnessed by the strengthened TAUTOLOGY gadget family).
@@
    Under the ETH (or when considering naive exhaustive search), exponential finding cost dominates linear maintenance cost for sufficiently large $n$.
Therefore, there exists $n_{0}$ such that for all $n>n_{0}$, over-modeling minimizes total expected cost:
@@
Over-modeling is economically optimal under the stated model and complexity constraints.
@@
By Theorem 3.6, SUFFICIENCY-CHECK is coNP-complete, so under the assumption $\textsf{P}\neq\textsf{coNP}$ there is no polynomial‑time algorithm that decides sufficiency for arbitrary inputs.
Finding the minimal sufficient set by brute force requires checking many candidate sets. Exhaustive search examines
$\sum_{i=0}^{n}\binom{n}{i}=2^{n}$
candidate subsets, so naive search has cost $O(2^{n})$. This shows only that brute-force is exponential. To assert a matching lower bound of $2^{\Omega(n)}$ for *all* algorithms requires additional assumptions. For example, under the Exponential Time Hypothesis (ETH), TAUTOLOGY (and hence SUFFICIENCY-CHECK in the succinct encoding) requires $2^{\Omega(n)}$ time, and SUFFICIENCY-CHECK inherits this conditional lower bound via our reduction.
@@
For concrete threshold: when $n=20$ parameters, exhaustive search requires $2^{20}\approx 10^{6}$ checks. Including $k=5$ extra parameters costs $O(5)$ maintenance overhead but avoids $10^{6}$ computational work.
Since $2^{n}$ grows faster than any polynomial in $k$ or $n$, there exists $n_{0}$ such that for all $n>n_{0}$:
@@
Adding underspecification risk $C_{\text{under}}$ (production failures from missing parameters), which is unbounded in the model, makes over-specification strictly dominant.  $\blacksquare$
@@
Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that:

### v3

@@ 5 Encoding-Regime Separation
Proposition 5.3 (Finite-State Query Lower-Bound Core via Empty-Set Subproblem).
In the finite-state query regime [Q_fin], let $S$ be any finite state type with $|S|\geq 2$ and let $Q\subset S$ with $|Q|<|S|$. Then there exist two decision problems $\mathcal{D}_{\mathrm{yes}},\mathcal{D}_{\mathrm{no}}$ over the same state space such that:
  • their oracle views on all states in $Q$ are identical,
  • $\emptyset$ is sufficient for $\mathcal{D}_{\mathrm{yes}}$,
  • $\emptyset$ is not sufficient for $\mathcal{D}_{\mathrm{no}}$.
Consequently, no deterministic query procedure using fewer than $|S|$ state queries can solve SUFFICIENCY-CHECK on all such instances; the worst-case query complexity is $\Omega(|S|)$. (L: CC50, L7, L13.)
Proof.
This is exactly the finite-state indistinguishability theorem CC4. For any $|Q|<|S|$, there is an unqueried hidden state $s_{0}$ producing oracle-indistinguishable yes/no instances with opposite truth values on the $I=\emptyset$ subproblem. Since SUFFICIENCY-CHECK contains that subproblem, fewer than $|S|$ queries cannot be correct on both instances, yielding the $\Omega(|S|)$ worst-case lower bound.  $\blacksquare$
Proposition 5.4 (Checking–Witnessing Duality on the Empty-Set Core).
[C:P5.4;R:H=query-lb] Let $n\geq 1$ and let $P$ be a finite family of Boolean-state pairs. Assume $P$ is sound for refuting empty-set sufficiency in the sense that for every $\text{Opt}:\{0,1\}^{n}\to\{0,1\}$, whenever $\emptyset$ is not sufficient, some pair in $P$ witnesses disagreement of Opt. Then
$|P|\geq 2^{n-1}.$
Equivalently, any sound checker must inspect at least the witness budget $W(n)=2^{n-1}$, so checking budget $T(n)$ satisfies $T(n)\geq W(n)$. (L: DQ587, DQ588.)
Corollary 5.5 (Information Barrier in Query Regimes).
[C:C5.5;R:H=query-lb] For the fixed sufficiency question, finite query interfaces induce an information barrier:
  • Q_fin
    (Opt-oracle): with $|Q|<|S|$, there exist yes/no instances indistinguishable on all queried states but with opposite truth values on the $I=\emptyset$ subproblem.
  • Q_bool
    value-entry and state-batch interfaces preserve the same obstruction at cardinality $2^{n}$.
Hence the barrier is representational-access based: without querying enough of the hidden state support, exact certification is blocked by indistinguishability rather than by search-strategy choice. (L: CC27, CC29, CC28.)
Corollary 5.6 (Boolean-Coordinate Instantiation).
[C:C5.6;R:H=query-lb] In the Boolean-coordinate state space $S=\{0,1\}^{n}$, Proposition 5.3 yields the familiar $\Omega(2^{n})$ worst-case query lower bound for Opt-oracle access. (L: CC49; core: L6.)
Proposition 5.7 (Value-Entry Query Lower Bound).
In the mechanized Boolean value-entry query submodel [Q_bool], for any deterministic procedure using fewer than $2^{n}$ value-entry queries $(a,s)\mapsto U(a,s)$, there exist two queried-value-indistinguishable instances with opposite truth values for SUFFICIENCY-CHECK on the $I=\emptyset$ subproblem. Therefore worst-case value-entry query complexity is also $\Omega(2^{n})$. (L: L9, L19.)
Proof.
The theorem L2 constructs, for any query set of cardinality $2^{n}$, an unqueried hidden state $s_{0}$ and a yes/no instance pair that agree on all queried values but disagree on $\emptyset$-sufficiency. The auxiliary bound L6 ensures that fewer than $2^{n}$ value-entry queries cannot cover all states, so the indistinguishability argument applies in the worst case.  $\blacksquare$
Proposition 5.8 (Subproblem-to-Full Transfer Rule).
If every full-problem solver induces a solver for a fixed subproblem, then any lower bound for that subproblem lifts to the full problem. (L: CC58, CC59, CC52.)
This is the restriction-map instance of Definition 2.23: a solver for the full regime induces one for the restricted subproblem regime, so lower bounds transfer.
Proposition 5.9 (Randomized Robustness (Seedwise)).
[C:P5.9;R:H=query-lb] In [Q_bool], for any query set with cardinality $2^{n}$, the indistinguishable yes/no pair from Proposition 5.3 forces one decoding error per random seed for any seed-indexed decoder from oracle transcripts. Consequently, finite-support randomization does not remove the obstruction: averaging preserves a constant error floor on the hard pair. (L: L11, L12, L5.)
Proposition 5.10 (Randomized Robustness (Weighted Form)).
[C:P5.10;R:H=query-lb] For any finite-support seed weighting $\mu$, the same hard pair satisfies a weighted identity: the weighted sum of yes-error and no-error equals total seed weight. Hence randomization cannot collapse both errors simultaneously. (L: L23, L24.)
Proposition 5.11 (State-Batch Query Lower Bound).
In [Q_bool], the same $\Omega(2^{n})$ lower bound holds for a state-batch oracle that returns the full Boolean-action utility tuple at each queried state. (L: L8, L14.)
Proposition 5.12 (Finite-State Empty-Subproblem Generalization).
[C:P5.12;R:H=query-lb] The empty-subproblem indistinguishability lower-bound core extends from Boolean-vector state spaces to any finite state type with at least two states. (L: L7, L13.)
Proposition 5.13 (Adversary-Family Tightness by Full Scan).
[C:P5.13;R:H=query-lb] For the const/spike adversary family used in the query lower bounds, querying all states distinguishes the pair; thus the lower-bound family is tight up to constant factors under full-state scan. (L: L10.)
Proposition 5.14 (Weighted Query-Cost Transfer).
[C:P5.14;R:H=query-lb] Let $w(q)$ be per-query cost and $w_{\min}$ a lower bound on all queried costs. Any cardinality lower bound $|Q|\geq L$ lifts to weighted cost:
$\sum_{q\in Q}w(q)\ \geq\ w_{\min}\cdot L.$
(L: L21, L22.)
Proposition 5.15 (Oracle-Lattice Transfer: Batch $\Rightarrow$ Entry).
In [Q_bool], agreement on state-batch oracle views over touched states implies agreement on all value-entry views for the corresponding entry-query set. (L: L20, CC43.)
This is the oracle-transducer instance of Definition 2.23: batch transcripts induce entry transcripts while preserving indistinguishability for the target sufficiency question.
Proposition 5.16 (Oracle-Lattice Strictness: Opt vs Value Entries).
‘Opt‘-oracle views are strictly coarser than value-entry views: there exist instances with identical ‘Opt‘ views but distinguishable value entries. (L: L3, L4.)
Remark 5.17 (Instantiation of Definitions 3.1 and 3.2).
Theorem 5.1 and Propositions 5.35.7 keep the structural problem fixed (same sufficiency relation) and separate representational hardness by access regime: explicit-state access exposes the boundary $s\mapsto\text{Opt}(s)$, finite-state query access already yields $\Omega(|S|)$ lower bounds at the Opt-oracle level (Boolean instantiation: $\Omega(2^{n})$), value-entry/state-batch interfaces preserve the obstruction in [Q_bool], and succinct access can hide structure enough to force ETH-level worst-case cost on a hard family.
This regime-typed landscape identifies tractability under [E], a finite-state query lower-bound core under [Q_fin] with Boolean interface refinements under [Q_bool], and worst-case intractability under [S+ETH] for the same underlying decision question.
The structural reason for this separation is enumerable access. Under the explicit-state encoding, the mapping $s\mapsto\text{Opt}(s)$ is directly inspectable and sufficiency verification reduces to scanning state pairs, at cost polynomial in $|S|$. Under succinct encoding, the circuit compresses an exponential state space into polynomial size; universal verification over that compressed domain, namely “for all $s,s^{\prime}$ with $s_{I}=s^{\prime}_{I}$, does $\text{Opt}(s)=\text{Opt}(s^{\prime})$?”, carries worst-case cost exponential in $n$, because the domain the circuit compressed away cannot be reconstructed without undoing the compression.
5.1 Budgeted Physical Crossover
The [E] vs [S] split can be typed against an explicit physical budget. Let $E(n)$ and $S(n)$ denote explicit and succinct representation sizes for the same decision question at scale parameter $n$, and let $B$ be a hard representation budget. These crossover claims are interpreted through the physical-to-core encoding discipline of Section 2.7: physical assumptions are transferred explicitly into core assumptions, and encoded physical counterexamples map to core failures on the encoded slice.[C:P2.34;C2.35;R:AR] (L: CT6,CT8,CT4.)
Proposition 5.18 (Budgeted Crossover Witness).
[C:P5.18;R:H=exp-lb-conditional] If $E(n)>B$ and $S(n)\leq B$ for some $n$, then explicit representation is infeasible at $(B,n)$ while succinct representation remains feasible at $(B,n)$. (L: CC45, PBC1.)
Proof.
This is exactly the definitional split encoded in PBC1: explicit infeasibility is $B<E(n)$ and succinct feasibility is $S(n)\leq B$. The theorem CC2 returns both conjuncts.  $\blacksquare$
Proposition 5.19 (Above-Cap Crossover Existence).
[C:P5.19;R:H=exp-lb-conditional] Assume there is a global succinct-size cap $C$ and explicit size is unbounded:
$\forall n,\ S(n)\leq C,\qquad\forall B,\ \exists n,\ B<E(n).$
Then for every budget $B\geq C$, there exists a crossover witness at budget $B$. (L: CC44, PBC2.)
Proof.
Fix $B\geq C$. By unboundedness of $E$, choose $n$ with $B<E(n)$. By the cap assumption, $S(n)\leq C\leq B$, so $(B,n)$ is a crossover witness. The mechanized closure theorem is CC1.  $\blacksquare$
Proposition 5.20 (Crossover Does Not Imply Exact-Certification Competence).
[C:P5.20;R:H=exp-lb-conditional] Fix any collapse schema “exact certification competence $\Rightarrow$ complexity collapse” for the target exact-certification predicate. Under the corresponding non-collapse assumption, exact-certification competence is impossible even when a budgeted crossover witness exists. (L: CC46.)
Proof.
The mechanized theorem CC3 bundles the crossover split with the same logical closure used in Proposition 3.61: representational feasibility and certification hardness are distinct predicates.  $\blacksquare$
Proposition 5.21 (Crossover Policy Closure Under Integrity).
[C:P5.21;R:H=exp-lb-conditional] In the certifying-solver model (Definitions 3.83.9), assume:
  • a budgeted crossover witness at $(B,n)$,
  • the same collapse implication as Proposition 3.61,
  • a solver that is integrity-preserving for the declared relation.
Then either full in-scope non-abstaining coverage fails, or declared budget compliance fails. Equivalently, in uncertified hardness-blocked regions, integrity is compatible with $\mathsf{ABSTAIN}/\mathsf{UNKNOWN}$ but not with unconditional exact-competence claims. (L: CC47.)
Corollary 5.22 (Finite Budget Implies Eventual Exact-Certification Failure).
[C:C5.22;R:H=exp-lb-conditional] Fix a physical computer model with finite total budget and strictly positive per-bit physical cost. If required operation count for exact certification has an exponential lower bound, then there exists a threshold $n_{0}$ such that for all $n\geq n_{0}$, required physical work exceeds budget. For the canonical requirement model used in this paper, the same eventual-budget-failure conclusion holds. (L: DQ569, DQ568, DQ574.)
Proposition 5.23 (Least Divergence Point at Fixed Budget).
[C:P5.23;R:H=exp-lb-conditional] Fix a budgeted encoding model and budget $B$. If a crossover witness exists at budget $B$, then there is a least index $n_{\mathrm{crit}}$ such that crossover holds at $n_{\mathrm{crit}}$, and no smaller index has crossover. (L: DQ576.)
Proposition 5.24 (Eventual Explicit Infeasibility from Monotone Growth).
[C:P5.24;R:H=exp-lb-conditional] If explicit size is monotone in $n$ and already exceeds budget $B$ at some $n_{0}$, then explicit infeasibility holds for all $n\geq n_{0}$. (L: DQ577.)
Proposition 5.25 (Payoff Threshold as Eventual Split).
[C:P5.25;R:H=exp-lb-conditional] If, beyond $n_{0}$, explicit encoding is infeasible while succinct encoding remains feasible, then every $n\geq n_{0}$ is a crossover point. (L: DQ575, DQ579.)
Corollary 5.26 (No Universal Survivor Without Succinct Bound).
[C:C5.26;R:H=exp-lb-conditional] If both explicit and succinct encodings are unbounded, then for every fixed budget $B$, there exists an index where explicit is infeasible and an index where succinct is infeasible. (L: DQ578.)
Proposition 5.27 (Policy Closure at and Beyond Divergence).
[C:P5.27;R:H=exp-lb-conditional] Under the same integrity-resource collapse assumption as Proposition 5.21, policy closure holds at any fixed divergence point and uniformly beyond any threshold where crossover holds. (L: DQ580, DQ581.)
5.2 Thermodynamic Lift in Declared Model
The crossover and hardness statements are complexity-theoretic. The thermodynamic lift follows the same assumption discipline as the rest of the paper: claims are exactly those derivable from the declared conversion model and declared bit-operation lower-bound premises.
Proposition 5.28 (Bit-to-Energy/Carbon Lift).
[C:P5.28;R:H=cost-growth] Fix a declared thermodynamic conversion model with constants $(\lambda,\rho)$ where $\lambda$ lower-bounds energy per irreversible bit operation and $\rho$ lower-bounds carbon per joule. If a bit-operation lower bound $b_{\mathrm{LB}}\leq b_{\mathrm{used}}$ holds, then:
$E_{\mathrm{LB}}=\lambda b_{\mathrm{LB}}\leq\lambda b_{\mathrm{used}},\qquad C_{\mathrm{LB}}=\rho E_{\mathrm{LB}}\leq\rho(\lambda b_{\mathrm{used}}).$
(L: CC65, CC66.)
Proposition 5.29 (Hardness–Thermodynamics Bundle).
[C:P5.29;R:H=cost-growth] Under the same non-collapse/collapse schema used by Proposition 3.61, exact-certification competence is impossible; combined with a declared bit-operation lower bound, this yields simultaneous lower bounds on energy and carbon in the declared conversion model. (L: CC67.)
Proposition 5.30 (Mandatory Physical Cost).
[C:P5.30;R:H=cost-growth] Within the declared linear thermodynamic model, if
$\lambda>0,\quad\rho>0,\quad b_{\mathrm{LB}}>0,$
then both lower bounds are strictly positive:
$E_{\mathrm{LB}}=\lambda b_{\mathrm{LB}}>0,\qquad C_{\mathrm{LB}}=\rho E_{\mathrm{LB}}>0.$
(L: CC68.)
Proposition 5.31 (Conserved Additive Accounting).
[C:P5.31;R:H=cost-growth] Within the same declared linear model, lower-bound accounting is additive over composed bit-operation lower bounds:
$E_{\mathrm{LB}}(b_{1}+b_{2})=E_{\mathrm{LB}}(b_{1})+E_{\mathrm{LB}}(b_{2}),\qquad C_{\mathrm{LB}}(b_{1}+b_{2})=C_{\mathrm{LB}}(b_{1})+C_{\mathrm{LB}}(b_{2}).$
(L: CC64.)
@@ 6 Tractable Special Cases: When You Can Solve It
We distinguish the encodings of Section 2.4. The tractability results below state the model assumption explicitly. Structural insight: hardness dissolves exactly when the full decision boundary $s\mapsto\text{Opt}(s)$ is recoverable in polynomial time from the input representation; the three cases below instantiate this single principle. Concretely, each tractable regime corresponds to a specific structural insight (explicit boundary exposure, additive separability, or tree factorization) that removes the hardness witness; this supports reading the general-case hardness as missing structural access in the current representation rather than as an intrinsic semantic barrier. In physical terms, these are the regimes where required structure is directly observable or factorized so that verification work scales with exposed local structure rather than hidden global combinations on encoded physical instances.[C:T6.1,2.37;P2.34;R:E,TA,AR] (L: CC69,CC70,CC71,CC72,CT6,CT4.)
@@ 6.4 Practical Implications
Proposition 6.2 (Heuristic Reusability).
[C:P6.2;R:H=tractable-structured] Let $\mathcal{C}$ be a structure class for which SUFFICIENCY-CHECK is polynomial (Theorem 6.1). If membership in $\mathcal{C}$ is decidable in polynomial time, then the combined detect-then-check procedure is a sound, polynomial-time heuristic applicable to all future instances in $\mathcal{C}$. For each subcase of Theorem 6.1, structure detection is polynomial (under the declared representation assumptions). (L: CC53, CC7, CC55, CC74.)
Remark 6.3 (Heuristics and Integrity).
Proposition 6.2 removes the complexity-theoretic obstruction to integrity-preserving action on detectable tractable instances: integrity no longer forces abstention. Whether the agent acts still requires competence (Definition 3.9): budget and coverage must also be satisfied. An agent that detects structure class $\mathcal{C}$, applies the corresponding polynomial checker, and abstains when detection fails maintains integrity; it never claims sufficiency without verification. Mistaking the heuristic for the general solution, claiming polynomial-time competence on a coNP-hard problem, violates integrity (Proposition 3.61).
7 Regime-Conditional Corollaries
This section derives regime-typed engineering corollaries from the core complexity theorems. Theorem 7.2 maps configuration simplification to SUFFICIENCY-CHECK; Theorems 4.6, 4.7, and 5.1 then yield exact minimization consequences under [S] and [S+ETH]. The same derivation applies to physical configuration/control systems once they are encoded into the C1–C4 decision interface: “configuration parameter” becomes any declared physical control or observation coordinate, and the same sufficiency predicate governs behavior preservation. [C:T7.2,2.37;P2.34;R:AR] (L: CR1,CT6,CT4.)
Regime tags used below follow Section 2.5: [S], [S+ETH], [E], [S_bool]. Any prescription that requires exact minimization is constrained by these theorem-level bounds.[C:T4.6,4.7,5.1;R:S,S+ETH] Theorem 7.4 implies that persistent failure to isolate a minimal sufficient set is a boundary-characterization signal in the current model, not a universal irreducibility claim.[C:T7.4;R:S]
Conditional rationality criterion.
For the objective “minimize verified total cost while preserving integrity,” persistent over-specification is admissible only in the attempted competence failure cell (Definition 3.59): after an attempted exact procedure, if exact irrelevance cannot be certified efficiently, integrity retains uncertified coordinates rather than excluding them.[C:P3.60;R:AR] When exact competence is available in the active regime (e.g., Theorem 6.1 and the exact-identifiability criterion), proven-irrelevant coordinates are removable; persistent over-specification is irrational for that same objective.[C:T6.1;P3.60;R:E,STR] Proposition 3.60 makes this explicit: in the integrity-preserving matrix, one cell is rational and three are irrational, so irrationality is the default verdict.[C:P3.60;R:AR] (L: IC18, IC25, IC7, CC71, S2P1.)
Remark 7.1 (Regime Contract for Engineering Corollaries).
All claims in this section are formal corollaries under the declared model assumptions.
  • Competence claims are indexed by the regime tuple of Definition 3.9; prescriptions are meaningful only relative to feasible resources under that regime (bounded-rationality feasibility discipline) [71, 5].[C:P3.58;R:AR]
  • Integrity (Definition 3.8) forbids overclaiming beyond certifiable outputs; $\mathsf{ABSTAIN}$$\mathsf{UNKNOWN}$ is first-class when certification is unavailable.[C:P3.58;R:AR]
  • Therefore, hardness results imply a regime-conditional trilemma: abstain, weaken guarantees (heuristics/approximation), or change encoding/structural assumptions to recover competence.[C:T4.6,5.1;P3.60;R:S,S+ETH]
@@ 7.1 Configuration Simplification is SUFFICIENCY-CHECK
Because SUFFICIENCY-CHECK is a meta-problem parameterized by an arbitrary decision problem $\mathcal{D}$, real engineering problems with factored configuration spaces are instances of the hardness landscape established above.
Theorem 7.2 (C1–C4: Configuration Simplification Reduces to SUFFICIENCY-CHECK).
Given a software system with configuration parameters $P=\{p_{1},\ldots,p_{n}\}$ and observed behaviors $B=\{b_{1},\ldots,b_{m}\}$, the problem of determining whether parameter subset $I\subseteq P$ preserves all behaviors is equivalent to SUFFICIENCY-CHECK. (L: CR1)
@@
  • Actions $A=B\cup\{\bot\}$, where $\bot$ is a sentinel “no-observed-behavior” action
@@
  • For $b\in B$, utility $U(b,s)=1$ if behavior $b$ occurs under configuration $s$, else 0
  • Sentinel utility $U(\bot,s)=1$ iff no behavior in $B$ occurs under configuration $s$, else 0
Then
[…]
So the optimizer map exactly encodes observed-behavior equivalence classes, including the empty-behavior case.
@@
Remark 7.3 (Reduction Scope).
The reduction above requires only:
  1. 1.
    a finite behavior set,
  2. 2.
    parameters with finite domains, and
  3. 3.
    an evaluable behavior map from configurations to achieved behaviors.
These are exactly the model-contract premises C1–C3 instantiated for configuration systems.
Theorem 7.4 (S: Over-Modeling as Boundary-Signal Corollary).
By contraposition of Theorem 7.2, if no coordinate set can be certified as exactly relevance-identifying (Definition 2.16) for the modeled system, then the decision boundary is not completely characterized by the current parameterization. (L: CC40, CC6)
Proof.
Assume the decision boundary were completely characterized by the current parameterization. Then, via Theorem 7.2, the corresponding sufficiency instance admits exact relevance membership, hence a coordinate set that satisfies Definition 2.16. Contraposition gives the claim: persistent failure of exact relevance identification signals incomplete characterization of decision relevance in the model.  $\blacksquare$
7.2 Cost Asymmetry Under ETH
We now prove a cost asymmetry result under the stated cost model and complexity constraints. [^1] Naive subset enumeration still gives an intuitive baseline of $O(2^{n})$ checks, but that is an algorithmic upper bound; the theorem below uses ETH for the lower-bound argument.
Theorem 7.5 (S+ETH: Cost-Asymmetry Consequence).
@@ 7.2 Cost Asymmetry Under ETH
Assume ETH in the succinct encoding model of Section 2.4. Then:
@@
    Exact identification of minimal sufficient sets has worst-case finding cost $C_{\text{find}}(n)=2^{\Omega(n)}$. (Under ETH, SUFFICIENCY-CHECK has a $2^{\Omega(n)}$ lower bound in the succinct model, and exact minimization subsumes this decision task.)
@@
    Under ETH, exponential finding cost dominates linear maintenance cost for sufficiently large $n$.
Therefore, there exists $n_{0}$ such that for all $n>n_{0}$, the finding-vs-maintenance asymmetry satisfies:
@@
Within [S+ETH], persistent over-specification is unresolved boundary characterization, not proof that all included parameters are intrinsically necessary. Conversely, when exact competence is available in the active regime, certifiably irrelevant coordinates are removable; persistence is irrational for the same cost-minimization objective.[C:T7.5;P3.60;R:S+ETH,TA] (L: CC12, HD13.)
@@
Under ETH, the TAUTOLOGY reduction used in Theorem 4.6 yields a $2^{\Omega(n)}$ worst-case lower bound for SUFFICIENCY-CHECK in the succinct encoding. Any exact algorithm that outputs a minimum sufficient set can decide whether the optimum size is 0 by checking whether the returned set is empty; this is exactly the SUFFICIENCY-CHECK query for $I=\emptyset$. Hence exact minimal-set finding inherits the same exponential worst-case lower bound.
@@
The eventual dominance step is mechanized in HD8: for fixed linear-overhead parameter $k$ and additive constant $C_{\text{under}}$ there is $n_{0}$ such that $k<2^{n}+C_{\text{under}}$ for all $n\geq n_{0}$. Therefore:
@@
For any fixed nonnegative $C_{\text{under}}$, the asymptotic dominance inequality remains and only shifts the finite threshold $n_{0}$.  $\blacksquare$
@@
[C:C7.6;R:H=succinct-hard] Assuming $\textsf{P}\neq\textsf{coNP}$, there exists no polynomial-time algorithm that: